40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Dec 8 15:59:21 2011, Frank, Summary, PMC, realignment 
    Reply  Thu Dec 8 20:28:06 2011, tara, Summary, PMC, realignment 
       Reply  Thu Dec 8 21:06:10 2011, Frank, Summary, PMC, realignment 
          Reply  Fri Dec 9 06:32:24 2011, Tara, Summary, PMC, realignment 
             Reply  Thu Jan 12 16:04:05 2012, Tara, Summary, PMC, Isolator base, EOM base's height faraday_base.jpg
                Reply  Fri Jan 13 16:51:29 2012, Tara, Notes, optic, Isolator base, EOM base's height eom_base_2011_01_13.PDFdual_periscope_base.PDF
                   Reply  Thu Jan 26 12:40:21 2012, Tara, Notes, optic, Dual periscope base dual_periscope_base.PDF
Message ID: 753     Entry time: Thu Dec 8 15:59:21 2011     Reply to this: 754
Author: Frank 
Type: Summary 
Category: PMC 
Subject: realignment 

started realigning everything from scratch and calibrate all channels right. PMC optical power channels need re-calibration.

input beam to PMC: 28.9mW
PMC transmitted beam: 23.4mw
transmission trough curved mirror: 1.63mW

--> only 81% visibility (87% including back mirror leakage), but should be enough for what we want to do. Don't know the loss mechanism (didn't investigate)

ELOG V3.1.3-